OpenStudy (anonymous):

A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week?

OpenStudy (anonymous):

You will have 4 times the second week amount of cookies, minus 5 (of the first week), in total

OpenStudy (anonymous):

so first week 104, second 109, third 218

OpenStudy (ash2326):

let the no. of boxes sold in first week x in second week x+5 in third week 2(x+5)=2x+10 total boxes sold=431 x+x+5+2x+10=431 4x=416 x=104 so in first week 104 boxes sold in second 109 in third 218

OpenStudy (anonymous):

thank u